PART 1: Find the present value for the following income stream if the interest rate is...

PART 1:

Find the present value for the following income stream if the interest rate is 7% percent and the payments occur at the end of each year. Cash flow for years 1-4: $750, years 5-11: $800, years 12-17: $1,100

A. $4,767

B. $8,321

C. $6,328

D. $5,189

PART 2:

How long will it take a sum of money to double in value if the annual interest rate is 10 percent? Round answer to one decimal.

A. 20.0 years
B. 7.3 years
C. 12.6 years
D. Insufficient information to determine – we need to know the amount of the initial investment
Answer #1

Part I:-

We will determine the Present Value using the PV Factors

For instance,

PV Factor for Year 1 = 1 / (1+7%) ^ 1 = 0.9345

PV Factor for Year 2 = 1 / (1+7%) ^ 2 = 0.8734

And so on..

Attaching the excel sheet with 2 images, one showing final value obtained and second showing formulas used:-

A. В C. E F 1 Years Cash Flows PV Factors Present Value of Cash Flows Discount Rate 7% 1 750 0.93458 700.9346 4 2 750 0.87344

A C E F 1 Present Value of Cash Flows Years Cash Flows Discount Rate 2 PV Factors | =1/(1+$F$3)^B3 =1/(1+$F$3)^B4 -1/(1+SF$3)

Hence, PV of Income Stream = $ 8,321 (Option B)

Part II :-

We will use the Excel function "NPER" to determine the number of years required to double the investment.

NPER function has the following syntax:-

NPER (Interest Rate, PMT, PV, FV)

Interest Rate = Annual Interest Rate of 10% in this problem

PMT = 0 (Since, no investment is required to be made yearly)

PV = Initial Investment made ( Let us assume that to be $100)

FV = 2*100 = $200

Attaching the excel snapshots :-

22 -100 23 PV 24 FV 200 10% 25 Interest Rate 26 NPER 7.2725409 27

23 PV -100 24 FV 200 25 Interest Rate 0.1 | =NPER(D25,0,D23,D24) 26 NPER 27

Hence, required no of years = 7.3 years (Option B)